A GAS STATION OWNER HAS 30 GALLONS OF GASOLINE WORTH P51.50 PER GALLON AND SOME WORTH P53.75 PER GALLON. HOW MANY GALLONS OF THE P53.75 BRAND MUST THE OWNER MIX IN ORDER TO PRODUCE A GASOLINE THAT COSTS P52 PER GALLON?
Answer by stanbon(75887)

A GAS STATION OWNER HAS 30 GALLONS OF GASOLINE WORTH P51.50 PER GALLON AND SOME WORTH P53.75 PER GALLON. HOW MANY GALLONS OF THE P53.75 BRAND MUST THE OWNER MIX IN ORDER TO PRODUCE A GASOLINE THAT COSTS P52 PER GALLON?
-------
equation:
value + value = value
53.75x + 51.5*30 = 52(x + 30)
-------
5375x + 5150*30 = 5200x + 5200*30
175x = 50*30
7x = 30
x = 2 6/7 gallons (amt. of 53.75 per gallon to add)
-------
Cheers,
Stan H.